OpenStudy (anonymous):

Find the 7th term of the sequence-1,-2,-4,-8

OpenStudy (amistre64):

can you determine a pattern?

OpenStudy (amistre64):

lets factor out the negative -(1,2,4,8,...) might make it clearer

OpenStudy (jiteshmeghwal9):

sorry its asking u to subtract the digit to itself to gt next digit becoz the digits are negative

OpenStudy (anonymous):

What do you mean?

OpenStudy (jiteshmeghwal9):

(-1-1)=(-2-2)=(-4-4) & so on to gt ur seventh term

OpenStudy (jiteshmeghwal9):

do u understand @lovejones2012 ???

OpenStudy (anonymous):

hmmmmmmmmm?

OpenStudy (jiteshmeghwal9):

(-4-4)=-8 (-8-8)=-16 (-9-9)=-18

OpenStudy (jiteshmeghwal9):

@amistre64 was also asking u to determine a pattern something like that:)

OpenStudy (amistre64):

-16-16 rather than -9-9 tho ;)

OpenStudy (jiteshmeghwal9):

ya!

OpenStudy (anonymous):

So the answer is -16- -16?

Parth (parthkohli):

Can't we just say that we are multiplying 2 to get the next term?

OpenStudy (anonymous):

I meant to say -16 -16??

Parth (parthkohli):

We are making this a little tedious by first subtracting 2 then 4 then 8.. this is a geometric sequence with -1 as \(a\) and 2 as \(r\)!

OpenStudy (amistre64):

if we call the "current term"; An if we call the "next term"; An+1 we can see that each An+1 term is equal to 2 times the An term

Parth (parthkohli):

\(T_n = T_{n - 1} \times r\)

OpenStudy (amistre64):

Parth, do the interaction of the asker indicate that they are that advanced yet?

Parth (parthkohli):

Sorry, Amistre.

OpenStudy (anonymous):

So what is the answer? I am confused.....

OpenStudy (amistre64):

the answer is what we get after we go thur the process

Parth (parthkohli):

To make it simple, you may just multiply the number that you have already been multiplying to get the next terms in the sequence.

Parth (parthkohli):

That special number is \(2\)!

Parth (parthkohli):

Not two factorial, but two.

OpenStudy (amistre64):

1st term = -1 2nd term = 1st term * 2 = -1*2 = -2 3rd term = 2nd term * 2 = -2*2 = -4 4th term = 3rd term * 2 = -4*2 = -8 and it follows this same pattern from 1st to 7th

OpenStudy (amistre64):

so, what would you say the 5th term would be if you follow the pattern laid out?

OpenStudy (anonymous):

?

OpenStudy (amistre64):

you will have to be clearer than that since i cannot read minds ....

OpenStudy (amistre64):

The intent of this website is not to hand out free answers; but to help you study the concepts so that you can better understand the material.